# LCLs beyond bounded degrees - Locally Finite Labelings

### v1

@@ 10 $C^{X}\leftarrow B$ $\triangleright$ update
Then we must add $d=m_{\ell}(C_{X})-m_{\ell}(C)=2$ additional required copies of optional neighbors. There are several valid choices: we may add two copies of $u$ (extending each by $T_{u}$), or two copies of $v$ (extending each by $T_{v}$), or one copy of each (extending by $T_{u}$ and $T_{v}$, respectively). Each choice yields a different configuration, but every resulting configuration has the property that any neighborhood matching it can also match both $C_{X}$ and $C$. See Figures 11 and 12 for an illustration of this toy example.
@@
Let $\mathcal{N}_{r}(v)$ be any labeled $r$-hop neighborhood that matches some $C^{\prime}\in C^{X}$, and let $f:\mathcal{N}_{r}(v)\to\mathcal{V}^{\prime}$ be a witnessing matching of $C^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathfrak{c}^{\prime},\mu^{\prime},\tau^{\prime})$. First, by Observation 18 and the same argument used in Lemma 15, the neighborhood $\mathcal{N}(v)$ also matches $C_{X}$.
It remains to show that $\mathcal{N}_{r}(v)$ matches $C=(\mathcal{V},\mathcal{E},\mathfrak{c},\mu,\tau)$. By Observation 18, every node of $C^{\prime}$ is a copy of some node of $C$, and the corresponding configbranches are isomorphic. Hence there is a natural surjective map $g:\mathcal{V}^{\prime}\to\mathcal{V}$ that sends each copy in $C^{\prime}$ to its original in $C$. Note that $g$ is a graph homomorphism (it preserves adjacency).
@@
    Required-subgraph isomorphism. View $C$ as a tree rooted at $\mathfrak{c}$. Let $e\in\mathcal{E}$ be any required edge of $C$, and let $u$ be the neighbor of $\mathfrak{c}$ that lies on the unique path from $\mathfrak{c}$ toward $e$. Since $e$ is required, the edge $\{\mathfrak{c},u\}$ is required as well (required edges form a connected subgraph containing the center). By Observation 18 there is a unique copy $u^{\prime}\in\mathcal{V}^{\prime}$ of $u$ in $C^{\prime}$, and the branch $T_{u^{\prime}}$ contains a copy $e^{\prime}\in\mathcal{E}^{\prime}$ of $e$ which is required. Because $f$ matches $C^{\prime}$, exactly one edge of $\mathcal{N}_{r}(v)$ is mapped to $e^{\prime}$. Call that edge $e^{*}$. Then $f^{\prime}(e^{*})=g(f(e^{*}))=g(e^{\prime})=e$. Moreover, uniqueness of the copy $u^{\prime}$ and the fact that $\{\mathfrak{c}^{\prime},u^{\prime}\}$ is required imply that $f^{\prime}$ is injective on the required subgraph; hence $f^{\prime}$ restricts to an isomorphism between the required subgraph of $\mathcal{N}_{r}(v)$ and that of $C$.
@@ 5.1 Computing types of trees
Lemmas 44 and 45 also hold in the unbounded-degree setting.
@@
The Compress procedure requires a distance-$\ell$ coloring with a constant number of colors. For bounded-degree graphs, this can be computed in $O(\log^{*}n)$ rounds. In the unbounded-degree setting such a coloring might not exist.
However, the nodes removed by Compress do not depend on this coloring; only the remaining nodes require it. We can therefore compute the distance coloring locally in the subgraph induced by the removed nodes, which consists only of long paths. Computing a distance coloring on such paths still takes $O(\log^{*}n)$ rounds.
Since the rest of the graph does not depend on this coloring, this results only in an additive $O(\log^{*}n)$ overhead, leaving the asymptotic complexities of Lemmas 44 and 45 unchanged. ∎
@@
Because all assigned types are non-empty, this propagation continues until leaf nodes are reached, yielding a correct solution to $\Pi$. As in 47, the total running time is $O(\gamma\cdot k)$ rounds, which yields $O(n^{1/k})$, if $f$ is $k$-good and $O(\log n)$, if $f$ is $\infty$-good. ∎
@@ 6 Outlook and Open Questions
Does there exist a correct notion of locally checkable problems in this lower regime for unbounded-degree trees? More specifically, which structural properties of a problem determine whether it can be solved efficiently when degrees are large?
@@
As discussed earlier, the complexity of MIS is $\Theta(\log n/\log\log n)$, or more precisely $\Theta(\min\{\Delta,\log_{\Delta}n\})$, showing that parts of the complexity landscape must necessarily change in the unbounded-degree setting, even for LFLs. We know that the problem of Sinkless Orientation remains $\Theta(\log n)$ even in the unbounded degree setting, whereas problems such as $(2,2)$-ruling sets—also encodable as LFLs—are known to have a $\sqrt{\Delta}$ dependence [5]. This motivates the following question.
@@
While the multiplicative $f(\Delta)$ dependencies in Theorem 64 are unavoidable for some problems, such as $\Delta$-ruling sets or distance-$\Delta$ coloring, these problems cannot be encoded as LFLs. They require referencing something at distance $\Delta$ and hence can only be encoded using pointer chains of length $\Delta$. These can only be realised with $\Omega(\Delta)$ distinct labels [^8] As long as the checking radius needs to be a constant.. In contrast, LFLs require a finite label set independent of the degree.
@@
Appendix B Complexities Beyond Square Root
In Section 3 we obtained polynomial complexities up to $\Theta(\sqrt{n})$ by enforcing paths with an increasing number of leaves. While elegant, this construction has a natural limitation: to distinguish nodes locally we must change the number of leaves attached, but doing so too frequently causes the total number of nodes required to grow so quickly that the resulting paths cannot be much longer than $\sqrt{n}$. Thus, it is difficult to push this approach beyond the square-root regime.
To overcome this barrier, we move from using simple leaf counts to attaching richer structures of radius two. This yields a vast combinatorial variety of possible attachments for a node, without requiring us to increase the number of nodes attached at each step. In practice, we achieve this through partition gadgets. A partition gadget provides many distinct local configurations of the same overall size, allowing us to make nodes locally distinguishable without rapidly consuming nodes. This, in turn, enables the construction of paths whose complexity grows strictly beyond $\Theta(\sqrt{n})$, albeit at the cost of additional complications in how we define the problem.
@@ Appendix B Complexities Beyond Square Root
$\mathcal{L}=\left\{\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{k}\right\}$
@@
The function $p(n)$ grows quickly, providing the desired combinatorial explosion, which we will exploit shortly. It is a classical result of Hardy and Ramanujan [18] that the partition function satisfies the asymptotic estimate:
@@
Let $\mathcal{L}=(\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{k})$ be a partition of some integer. To construct the $\mathcal{L}$-gadget attached to a node $v$, we proceed as follows. For each $1\leq i\leq k$:
@@
    if $\mathcal{L}_{i}=1$, attach a single leaf node directly to $v$;
@@
    if $\mathcal{L}_{i}>1$, attach a new node $u_{i}$ to $v$, and then attach $\mathcal{L}_{i}-1$ leaf nodes to $u_{i}$.
In other words, each part $\mathcal{L}_{i}$ of the partition corresponds either to a single leaf or to a small star centered at a new node $u_{i}$. Distinct partitions thus lead to structurally different neighborhoods of radius $2$, while keeping the overall gadget size bounded. This gives us the desired local variety without inflating the node count too quickly.
Next, we explain how to combine partition gadgets into paths. This construction is directly analogous to the $\alpha$-paths from Section 3, but instead of increasing the number of leaves linearly, we move through a sequence of distinct partition gadgets.
Let $0<\alpha<1$ and fix a natural number $x$. Define the gadget size as $g=\lfloor x^{\alpha}\rfloor$. Enumerate all partitions of $g$ in some arbitrary order $(\mathcal{L}_{i})_{i\in[p(g)]}$. A gadget path is then a path whose nodes $v_{i}$ carry these gadgets in sequence. Because $p(g)$ is exponentially large in $\sqrt{g}$, we have more than enough gadgets available to sustain long paths. However, to keep the total number of nodes under control, we truncate the path after $\tfrac{1}{2}\lfloor x^{1-\alpha}\rfloor$ gadgets.
@@
Let $0<\alpha<1$ be any real number and $x$ any natural number, such that $g:=x^{\alpha}$ is also a natural number, a path
@@
is an $(x,\alpha)$ gadget path if each $v_{i}$ has the partition gadget $\mathcal{L}_{i}$ attached (for some fixed ordering of all partitions of $g$) and
$1<L\leq\tfrac{1}{2}\lfloor x^{1-\alpha}\rfloor.$
The key property of this definition is that despite the enormous variety of gadgets, the total number of nodes remains bounded by $x$. In any graph on $n$ nodes, the length of any $(x,\alpha)$ gadget path is restricted by (i) the inherent length restriction of the definition (ii) the fact that for large values of $g$ the graph construction uses a lot of nodes. We obtain the following corollary.
@@
An $(x,\alpha)$-gadget path of length $L$ contains exactly
$L+L\cdot x^{\alpha}$
nodes. Consequently, for any fixed $0<\alpha<1$ and all allowed  [^9] Here allowed is with respect to the definition of gadget paths, that is $x^{\alpha}$ must be an integer. integers $x$, the length of the longest $(x,\alpha)$-gadget path in an $n$-node graph is bounded by $O(n^{1-\alpha})$.
@@
We are now ready to define the actual LCL problem. The idea is exactly analogous to the $\alpha$-path construction: a node must output $1$ if and only if it belongs to a valid gadget path, and must output 0 otherwise. To make this precise, we must incorporate the notion of gadget paths into the problem definition for every relevant gadget size $g$.
@@
For each integer $x\geq 1$, such that $g:=x^{\alpha}$ is also a natural number, let $(\mathcal{L}_{i})_{i\in[p(g)]}$ be a fixed ordering of all partitions of $g$, and let $(x,\alpha)$ gadget paths be defined as above.
An instance of $\Pi^{\mathrm{part}}$ is a tree $T=(V,E)$ together with labels $\{0,1\}$ on the nodes. The labeling is locally correct if and only if every node $v\in V$ satisfies the following:
  • If $v$ is contained in a valid gadget path, then $v$ must output $1$.
@@
This definition generalizes the previous $\alpha$-path problems: instead of leaf counts alone, we now allow richer partition gadgets as the local structures that enforce long paths.
Observation 71.
@@
    The problem is 33 checkable, as defined at the beginning of Section 3.
@@
    The problem can be represented in local configurations as in Section 3.
@@
First, each node $v$ can verify whether or not it has a gadget attached using information from within distance 2. From this information, it can also determine precisely which gadget is attached and therefore deduce the corresponding parameter $x$. Moreover, to identify the gadgets attached to its neighbors, $v$ requires information up to distance 3. As in Section 3 (and as will be illustrated below), this suffices to represent the problem in terms of local configurations.
Second, to represent the problem as local configurations, we need the same basic set of configurations as before, but now one such set (with infinitely many configurations) for each possible choice of $x$. This generic construction is illustrated in Figure 15. Importantly, the arbitrary extensions of these configurations (i.e., the dotted nodes marked “arbitrary non-gadget neighbors”) must be constrained so that they cannot themselves form part of a gadget.
To see why this is necessary, consider the top configuration in Figure 15, and assume that the center node $v$ (marked by the arrow) has a gadget $G_{x}(\mathcal{L}_{i})$ attached. Suppose further that $v$ also has a neighbor $u$ which itself has only nine degree-one neighbors (i.e., leaves). In this case, $v$ would not actually have gadget $G_{x}(\mathcal{L}_{i})$ attached; instead, the combined structure would correspond to a partition of a number that is larger by ten. To avoid such ambiguities, all neighbors of $v$ that are either leaves themselves or that have only leaves (other than $v$) are considered part of a single gadget.
Figure 15: Key local configurations of $\Pi(\alpha)$, with center nodes indicated by arrows. Nodes without an explicitly specified output may take an arbitrary label (0 or 1). The $G_{x}(\mathcal{L}_{i})$ cones indicate that gadget $\mathcal{L}_{i}$ of the $(\alpha,x)$ gadget path is attached to a node.
     Bottom : Start of an $(x,\alpha)$ gadget path. The center node must output 1 if it has gadget $G_{x}(\mathcal{L}_{1})$ attached and it has a neighbor with the $G_{x}(\mathcal{L}_{2})$ gadget attached.
     Top: Propagation within an $(x,\alpha)$ gadget path. The center node has gadget $G_{x}(\mathcal{L}_{i})$ attached (for some $i\in\mathbb{N}$), and one of its neighbors must have gadget $G_{x}(\mathcal{L}_{i-1})$ attached and output 1. So the configuration essentially requires a predecessor. In the same way as before, this configuration also acts as an end to an $(x,\alpha)$ gadget path.
Now the only thing left to do, is to go through the same argument as in Section 3 to obtain all complexities See 1
@@
Upper bound. Consider the deterministic algorithm where each node $v$ first determines the largest partition gadget attached to it. By counting the size $g$ of this gadget, it can deduce the corresponding parameter $x=g^{1/\alpha}$, and hence identify its exact position in a valid $(x,\alpha)$-gadget path. By Corollary 69, it is sufficient to explore a neighborhood of radius $O(n^{1-\alpha})$ to decide whether $v$ belongs to such a path. Therefore, after at most $O(n^{1-\alpha})$ rounds in the LOCAL model, node $v$ can correctly output $1$ if it lies on a valid gadget path, and 0 otherwise. By construction, this yields a correct solution.
@@
Assume, for contradiction, that there exists a constant $\varepsilon>0$ such that $T(n)\leq\varepsilon n^{1-\alpha}$ and the algorithm succeeds with probability greater than $1/2$. Choose $n$ large enough so that these conditions hold and additionally ensure that $n^{\alpha}$ is an integer. Consider an $(n,\alpha)$-gadget path
@@
of length $L=\tfrac{1}{2}n^{1-\alpha}$, which by Corollary 69 uses fewer than $n$ nodes. We can extend $P$ to an $n$-node tree by, for example, attaching an additional path to $v_{L}$, which preserves the gadget-path property.
@@
In $G_{1}$, the node $v_{L}$ must output $1$, while in $G_{2}$, the same node must output 0. However, since $L=\tfrac{1}{2}n^{1-\alpha}>\varepsilon n^{1-\alpha}\geq T(n)$, the $T(n)$-neighborhood of $v_{L}$ is identical in both graphs. Thus $v_{L}$ follows the same output distribution in $G_{1}$ and $G_{2}$, and therefore with probability at least $1/2$ the algorithm errs on one of the two instances.

### v2

@@ 10 $C^{X}\leftarrow B$ $\triangleright$ update
Then we must add $d=m_{\ell}(C_{X})-m_{\ell}(C)=2$ additional required copies of optional neighbors. There are several valid choices: we may add two copies of $u$ (extending each by $T_{u}$), or two copies of $v$ (extending each by $T_{v}$), or one copy of each (extending by $T_{u}$ and $T_{v}$, respectively). Each choice yields a different configuration, but every resulting configuration has the property that any neighborhood matching it can also match both $C_{X}$ and $C$. See Figures 10 and 11 for an illustration of this toy example.
@@
Let $\mathcal{N}_{r}(v)$ be any labeled $r$-hop neighborhood that matches some $C^{\prime}\in C^{X}$, and let $f:\mathcal{N}_{r}(v)\to\mathcal{V}^{\prime}$ be a witnessing matching of $C^{\prime}=(\mathcal{V}^{\prime},\mathcal{E}^{\prime},\mathfrak{c}^{\prime},\mu^{\prime},\tau^{\prime})$. First, by Observation 14 and the same argument used in Lemma 11, the neighborhood $\mathcal{N}(v)$ also matches $C_{X}$.
It remains to show that $\mathcal{N}_{r}(v)$ matches $C=(\mathcal{V},\mathcal{E},\mathfrak{c},\mu,\tau)$. By Observation 14, every node of $C^{\prime}$ is a copy of some node of $C$, and the corresponding configbranches are isomorphic. Hence there is a natural surjective map $g:\mathcal{V}^{\prime}\to\mathcal{V}$ that sends each copy in $C^{\prime}$ to its original in $C$. Note that $g$ is a graph homomorphism (it preserves adjacency).
@@
    Required-subgraph isomorphism. View $C$ as a tree rooted at $\mathfrak{c}$. Let $e\in\mathcal{E}$ be any required edge of $C$, and let $u$ be the neighbor of $\mathfrak{c}$ that lies on the unique path from $\mathfrak{c}$ toward $e$. Since $e$ is required, the edge $\{\mathfrak{c},u\}$ is required as well (required edges form a connected subgraph containing the center). By Observation 14 there is a unique copy $u^{\prime}\in\mathcal{V}^{\prime}$ of $u$ in $C^{\prime}$, and the branch $T_{u^{\prime}}$ contains a copy $e^{\prime}\in\mathcal{E}^{\prime}$ of $e$ which is required. Because $f$ matches $C^{\prime}$, exactly one edge of $\mathcal{N}_{r}(v)$ is mapped to $e^{\prime}$. Call that edge $e^{*}$. Then $f^{\prime}(e^{*})=g(f(e^{*}))=g(e^{\prime})=e$. Moreover, uniqueness of the copy $u^{\prime}$ and the fact that $\{\mathfrak{c}^{\prime},u^{\prime}\}$ is required imply that $f^{\prime}$ is injective on the required subgraph; hence $f^{\prime}$ restricts to an isomorphism between the required subgraph of $\mathcal{N}_{r}(v)$ and that of $C$.
@@ 4.1 Computing Types of Trees
Lemmas 40 and 41 also hold in the unbounded-degree setting.
@@
The Compress procedure requires a distance-$\ell$ colouring with a constant number of colours. For bounded-degree graphs, this can be computed in $O(\log^{*}n)$ rounds. In the unbounded-degree setting such a colouring might not exist.
However, the nodes removed by Compress do not depend on this colouring; only the remaining nodes require it. We can therefore compute the distance colouring locally in the subgraph induced by the removed nodes, which consists only of long paths. Computing a distance colouring on such paths still takes $O(\log^{*}n)$ rounds.
Since the rest of the graph does not depend on this colouring, this results only in an additive $O(\log^{*}n)$ overhead, leaving the asymptotic complexities of Lemmas 40 and 41 unchanged. ∎
@@
Because all assigned types are non-empty, this propagation continues until leaf nodes are reached, yielding a correct solution to $\Pi$. As in 43, the total running time is $O(\gamma\cdot k)$ rounds, which yields $O(n^{1/k})$, if $f$ is $k$-good and $O(\log n)$, if $f$ is $\infty$-good. ∎
@@ 5 Outlook and Open Questions
More specifically, which structural properties of a problem determine whether it can be solved efficiently when degrees are large?
@@
As discussed earlier, the complexity of MIS is $\Theta(\log n/\log\log n)$, or more precisely $\Theta(\min\{\Delta,\log_{\Delta}n\})$, showing that the lower regime of the complexity landscape must necessarily change in the unbounded-degree setting, even for LFLs. We know that the problem of Sinkless Orientation remains $\Theta(\log n)$ even in the unbounded degree setting, whereas problems such as $(2,2)$-ruling sets—also encodable as LFLs—are known to have a $\sqrt{\Delta}$ dependence [5]. This motivates the following question.
@@
While the multiplicative $f(\Delta)$ dependencies in Theorem 60 are unavoidable for some problems, such as $\Delta$-ruling sets or distance-$\Delta$ colouring, these problems cannot be encoded as LFLs. They require referencing something at distance $\Delta$ and hence can only be encoded using pointer chains of length $\Delta$. These can only be realised with $\Omega(\Delta)$ distinct labels [^9] As long as the checking radius needs to be a constant.. In contrast, LFLs require a finite label set independent of the degree.
@@
The above result by [13] already gives us an idea of what the correct structural property is for problems that are $o(\log_{\Delta}n)$. Namely that these problems admit fast SLOCAL algorithms.
@@
Appendix B LCLs with Infintely Many Configurations
We present a formal version of the construction described in the introduction. For this construction, we retain the Local of Locally Checkable Loabelings in the sense that solutions depend only on the local information around a node. To this end, we call a problem $r$-checkable if
  1. 1.
  2. 2.
While in the introduction we enumerated all possible rooted trees with a certain number of nodes, we do the same here formally by using partitions of integers. In principle each such partition represents a possible height 2 tree, which we will call a partition gadget.
These partition gadgets provide many distinct local configurations of the same overall size, allowing us to make nodes locally distinguishable without rapidly consuming nodes.
@@ Appendix B LCLs with Infintely Many Configurations
$\mathcal{L}=\left\{a_{1},a_{2},\ldots,a_{k}\right\}$
@@
The function $p(n)$ grows quickly, providing the combinatorial explosion that we will exploit shortly. It is a classical result of Hardy and Ramanujan [18] that the partition function satisfies the asymptotic estimate:
@@
Let $\mathcal{L}=(a_{1},a_{2},\ldots,a_{k})$ be a partition of some integer. To construct the $\mathcal{L}$-gadget attached to a node $v$, we proceed as follows. For each $1\leq i\leq k$:
@@
    if $a_{i}=1$, attach a single leaf node directly to $v$;
@@
    if $a_{i}>1$, attach a new node $u_{i}$ to $v$, and then attach $a_{i}-1$ leaf nodes to $u_{i}$.
In other words, each part $a_{i}$ of the partition corresponds either to a single leaf or to a small star centered at a new node $u_{i}$. Distinct partitions thus lead to structurally different neighborhoods of radius $2$, while keeping the overall gadget size bounded.
Next, we explain how to combine partition gadgets into paths.
Let $0<\alpha<1$ and fix a natural number $g$. We fix an ordering $(\mathcal{L}_{i})_{i\in[p(g)]}$ of all partitions of $g$ for each possible value of $g$. A gadget path is then a path whose nodes $v_{i}$ carry these gadgets in sequence. Because $p(g)$ is exponentially large in $\sqrt{g}$, we have more than enough gadgets available to sustain long paths. However, to keep the total number of nodes under control, we truncate the path after $\tfrac{1}{2}\lfloor g^{\frac{1}{\alpha}-1}\rfloor$ gadgets.
If we choose $g$ large enough with respect to $\alpha$, then we always have enough gadgets. The following is an immediate consequence of Lemma 62.
Corollary 64.
Let $0<\alpha<1$, there exists a constant $g_{\alpha}$, such that all integers $g>g_{\alpha}$ satisfy
$g^{\frac{1}{\alpha}-1}<p(g)$
Now that we know that we have enough gadgets, we can define our length constrained gadget paths.
@@
Fix a unique order $(\mathcal{L}_{i})_{i\in[p(n)]}$ for the partitions of each natural number $n$.
Let $0<\alpha<1$ be any real number and let $g$ be any integer large enough s.t.
$g^{\frac{1}{\alpha}-1}<p(g)$
then a path
@@
in any graph is an $(\alpha,g)$ gadget path if each $v_{i}$ has the $i$-th partition gadget for $g$ attached and the length $L$ is bounded by
$1<L\leq\tfrac{1}{2}g^{\frac{1}{\alpha}-1}$
The key property of this definition is that despite the enormous variety of gadgets, the total number of nodes remains bounded with respect to $g$. In any graph on $n$ nodes, the length of any $(\alpha,g)$ gadget path is restricted by (i) the inherent length restriction of the definition (ii) the fact that for large values of $g$ the graph construction uses a lot of nodes. We obtain the following corollary.
@@
An $(\alpha,g)$-gadget path of length $L$ contains exactly
$L+L\cdot g$
nodes. Consequently for any fixed $0<\alpha<1$ and all allowed integers $g$, the length of the longest $(\alpha,g)$-gadget path in an $n$-node graph is in $\Theta(n^{1-\alpha})$.
The second part comes from the fact that, by definition, the lengths $L$ of the gadget paths are bounded by $\tfrac{1}{2}g^{\frac{1}{\alpha}-1}$.
@@
We are now ready to define the actual locally checkable problem: a node must output $1$ if and only if it belongs to a valid gadget path, and must output 0 otherwise.
@@
An instance of $\Pi(\alpha)$ is a tree $T=(V,E)$ where nodes are required to output one of $\{0,1\}$. The labeling is locally correct if and only if every node $v\in V$ satisfies the following:
  • If $v$ is contained in a $(\alpha,g)$-gadget path for some valid $g$, then $v$ must output $1$.
@@
we make the following observations:
Observation 68.
@@
    The problem is $4$ checkable.
@@
    The problem can be represented in local configurations.
@@
First, each node $v$ can verify whether or not it has a gadget attached using information from within distance 3. The nodes of $v$’s gadget are at distance 2 from $v$, so checking whether these nodes are leafes requires distance 3 information [^10] Refer to the definitions in the preliminaries to verify exactly how these distances are defined.. Moreover, to identify the gadgets attached to its neighbors, $v$ requires information up to distance 4. From this information, $v$ can determine the corresponding parameter $g$ and which gadget is attached to itself and its neighbor.
Second, to represent the problem as local configurations, we need a configuration that indicates the start of such a path and one which handles the propagation along the path. We require these configuraitons for every possible choice of $g$. This generic construction is illustrated in Figure 14. Importantly, the arbitrary extensions of these configurations (i.e., the dotted nodes marked “arbitrary non-gadget neighbors”) must be constrained so that they cannot themselves form part of a gadget.
To see why this is necessary, consider the top configuration in Figure 14, and assume that the center node $v$ (marked by the arrow) has a gadget $G_{g}(\mathcal{L}_{i})$ attached. Suppose further that $v$ also has a neighbor $u$ which itself has only nine degree-one neighbors (i.e., leaves). In this case, $v$ would not actually have gadget $G_{g}(\mathcal{L}_{i})$ attached; instead, the combined structure would correspond to a partition of a number that is larger by ten. To avoid such ambiguities, all neighbors of $v$ that are either leaves themselves or that have only leaves (other than $v$) are considered part of a single gadget.
Figure 14: Key local configurations of $\Pi(\alpha)$, with center nodes indicated by arrows. Nodes without an explicitly specified output may take an arbitrary label (0 or 1). The $G_{g}(i)$ cones indicate that the gadget of the $i$-th partition of $g$ is attached to a node.
     Bottom : Start of an $(\alpha,g)$ gadget path. The center node must output 1 if it has gadget $G_{g}(1)$ attached and it has a neighbor with the $G_{g}(2)$ gadget attached.
     Top: Propagation within an $(\alpha,g)$ gadget path. The center node has gadget $G_{g}(i)$ attached (for some $i\in\mathbb{N}$), and one of its neighbors must have gadget $G_{g}(i-1)$ attached and output 1. So the configuration essentially requires a predecessor. This configuration only exists up until $i\leq\frac{1}{2}g^{\frac{1}{\alpha}-1}$ for each $g$.
Now the only thing left to do, is proof that our construction allows us to obtain any polynomial complexity. See 1
@@
Upper bound. Consider the deterministic algorithm where each node $v$ first determines the largest partition gadget attached to it. By counting the size $g$ of this gadget, it can determine the $(\alpha,g)$-path constraints that it needs to satisfy (and exactly where its attached gadget lies in the fixed ordering of all partitions of $g$). By Corollary 66, it is sufficient to explore a neighborhood of radius $O(n^{1-\alpha})$ to decide whether $v$ belongs to such a path. Therefore, after at most $O(n^{1-\alpha})$ rounds in the LOCAL model, node $v$ can correctly output $1$ if it lies on a valid gadget path, and 0 otherwise. By construction, this yields a correct solution.
@@
Assume, for contradiction, that there exists a constant $\varepsilon>0$ such that $T(n)\leq\varepsilon n^{1-\alpha}$ and the algorithm succeeds with probability greater than $1/2$. Choose $n$ large enough so that these conditions hold and such that $g:=\lfloor n^{\alpha}\rfloor$ is large enough for Corollary 64.
Consider an $(\alpha,g)$-gadget path
@@
of length
$L=\tfrac{1}{4}\lfloor n^{1-\alpha}\rfloor<\frac{1}{2}g^{\frac{1}{\alpha}-1}$
, which by Corollary 66 uses fewer than $n$ nodes. We can extend $P$ to an $n$-node tree by, for example, attaching an additional path to $v_{L}$, which preserves the gadget-path property.
@@
In $G_{1}$, the node $v_{L}$ must output $1$, while in $G_{2}$, the same node must output 0. However, since
$L=\tfrac{1}{4}\lfloor n^{1-\alpha}\rfloor>\varepsilon n^{1-\alpha}\geq T(n)$
, the $T(n)$-neighborhood of $v_{L}$ is identical in both graphs. Thus $v_{L}$ follows the same output distribution in $G_{1}$ and $G_{2}$, and therefore with probability at least $1/2$ the algorithm errs on one of the two instances.